y^3-2y^2+3y-4
y=5
can you show me the steps to get the right answer

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

y^3-2y^2+3y-4
y=5
----
5^3 - 2*5^2 + 3*5 - 4
----
= 125 - 2*25 + 15 - 4
----
= 75 + 15 -4
---
= 90-4
-----
= 86
============
Cheers,
Stan H.
==============

Answer by solver91311(24713)



In order to develop an answer, right or otherwise, one has to begin with a question. Since you did not ask a question other than "can you show me the steps to get the right answer?", question mark assumed, I have to ask: The right answer to what?

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism